Circular No. 5888

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993ae IN UGC 1071 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova of mag V = 16.5 on two films taken Nov. 7.92 and 7.97 UT. The object is located at α = 1^{h}27^{m}16^{s}.08, δ = −2^{o}14′05′′.5 (equinox 1950.0); the galaxy’s center is located at end figures 15^{s}.00, 29′′.9. Nothing was found at this position to mag 21.5 on films taken Oct. 20.07 and 20.11. A nearby star (V ∼ 14) is at end figures 15^{s}.42, 15′24′′.6.

E. Cappellaro, European Southern Observatory; and A. Bragaglia, Osservatorio Astronomico di Bologna, report: “On Nov. 11.2 UT, we obtained a spectrogram (range 340-910 nm, resolution 7 nm) of SN 1993ae, using the ESO 1.5-m telescope and the Boller & Chivens spectrograph. The spectrum is very similar to that of a normal type-Ia supernova ∼ 10 days after maximum. In particular, strong absorptions are measured at 386.2, 452.4, 519.5, 580.8, and 626.5 nm. The last of these, attributed to Si II 635.5 nm, indicates an expansion velocity of ∼ 10 000 km/s.”

X-RAY NOVA IN VELA Y. Tanaka, Institute of Space and Astronautical Science, on behalf

of the ASCA Team, communicates: “GRS 1009–45 (IAUC 5864, 5886) was observed with ASCA on Nov. 10 UT. The source was found in the BATSE error box, and its location is determined as α = 10^{h}11^{m}26^{s}.2, δ = −44^{o}48′23′′ (equinox 1950.0; systematic error-circle radius 1′). The spectrum is very soft (kT ∼ 0.6 keV if approximated by a thermal spectrum) and accompanied by a hard tail, which is very similar to GS 2000+25 and GS/GRS 1124–68. The source is still very bright, with a flux of ∼ 2 × 10^{−8} erg cm^{−2} s^{−1} or 0.8 Crab in the range 1-10 keV, of which the flux below 3 keV dominates.”

SUPERNOVA 1993ac IN ANONYMOUS GALAXY D. D. Balam reports a position for this object, obtained with the 0.5-m

reflector (+ CCD) at Climenhaga Observatory (University of Victoria): α = 5^{h}41^{m}36^{s}.99, δ = +63^{o}20′57′′.2 (equinox 1950.0; mean error 0′′.5). The nucleus of the host galaxy is at end figures 38^{s}.04, 24′′.3.

1993 November 12 Daniel W. E. Green